# On the Verification of Control Flow Attestation Evidence

### v1

@@ 3.1 Evidence Types
Figure 2: Illustration of logging a conditional branch (CB), indirect/direct calls (IC/DC), and returns (RET) in $\mathcal{C}F_{Log}$ with (a) hash-based, (b) verbatim, and (c) hybrid evidence.
3.1.1 Hash-based Evidence (E1)
@@
As control flow transfers occur, branch destination addresses are accumulated into a single hash chain. This is performed by first accumulating the current destination address with the previous hash and then hashing the accumulated result [4, 62, 56, 52], as exemplified in Fig. 2.a. After the first branch occurs, its destination 0xe004 is accumulated into the hash. Since this is the first branch destination, it is accumulated with an initial constant to obtain $H_{0}$. Upon the second branch instruction, its destination address 0xf000 is accumulated with $H_{0}$ to obtain $H_{1}$. This process continues until the end of the requested execution generating a unique value that represents the path. Although simple control flows produce a single hash in E1, programs with more complex control flows (including those with nested loops and branching) might require a set of hash digests [4, 52]. This approach yields minimal storage/transmission overheads (i.e., the fixed size of a hash digest). However, $\sf{\mathcal{V}rf}$ must perform its analysis solely based on the final hash digest and the executable binary.
3.1.2 Verbatim Evidence (E2)
@@
On the other side of the spectrum, more recent CFA techniques advocate for storing and transmitting $\mathcal{C}F_{Log}$ in its entirety [49, 51, 53, 6, 54, 10]. As depicted in Fig. 2.b, the full sequence of destination addresses is maintained in a buffer. Therefore, this approach requires further (lossless) compression or is limited to small and self-contained operations to avoid filling the dedicated storage too quickly. A common strategy is to log simple loops (those without internal branches) with their address once and a count/condition pertaining to the number of iterations performed [6, 49, 54, 10]. Another reduction is obtained by recording statically defined branch destinations as a single bit [51]. Since static destinations can be determined from the binary alone, branch instructions with statically defined destinations can be recorded as a single bit (1/0) to determine if the branch was taken or not. Branches with dynamically defined destinations (e.g., returns and indirect calls) are still recorded with their full addresses. If $\mathcal{C}F_{Log}$ remains larger than available storage despite these optimizations, it can be split into a set of multiple fixed-size slices and streamed to $\sf{\mathcal{V}rf}$, as suggested in [53, 6, 10]. Despite the aforementioned challenges, E2 provides $\sf{\mathcal{V}rf}$ with full control flow path information for analysis.
3.1.3 Hybrid Evidence (E3)
@@
Some techniques implement a combination of E1 and E2. OAT [50] and ARI [63] record forward edges verbatim and accumulate backward edges (returns) into a single hash chain. As shown in Fig. 2.c, the RoT in $\sf{\mathcal{P}rv}$ maintains a trace of forward edges ($T_{F}$) and a hash chain of returns ($H_{R}$). Since 0xf000 is the destination of an indirect call, its full address must be recorded whereas the first destination can be represented by a single bit ‘0’ encoding the ‘not-taken’ conditional branch address. After the attested execution ends, $T_{F}$ and $H_{R}$ are sent to $\sf{\mathcal{V}rf}$.
3.2 $\sf{\mathcal{V}rf}$’s Perspective: Detection & Remediation from E1, E2 & E3
@@
Consider an attack detection strategy in which $\sf{\mathcal{V}rf}$ constructs the CFG of the executable binary, i.e., as suggested in [4, 6, 62, 52, 53, 56, 54]. An initial set of invalid paths is defined as forward edges that do not exist in the statically constructed CFG. As valid indirect targets (e.g., return addresses or indirect calls) are over-approximated using the static CFG [64], $\sf{\mathcal{V}rf}$ may also use $\mathcal{C}F_{Log}$ along with the program’s binary to emulate the execution of the program locally. In this process, a Shadow Stack (SS) can be emulated locally to verify backward edges (return destinations) and data flow analysis can be employed to determine a set of valid call sites for each indirect call, updating the CFG accordingly. We observe feasibility of this verification strategy varies greatly depending on evidence types E1, E2, and E3.
Figure 3: SABRE Workflow
@@
For E1, with the only evidence available being a hash digest, $\sf{\mathcal{V}rf}$ must infer which path has resulted in the digest. To that end, it must generate each possible legal path based on the CFG and use an emulated SS to accumulate the path’s hash during each step of the CFG traversal. However, the complexity of this task grows exponentially with the number of branches, leading to the intractable path explosion problem [65, 66]. Considering anomalous/illegal paths, the analysis is even more challenging. $\sf{\mathcal{V}rf}$ would need to map the hash digest to the exact illegal path. Illegal paths do not exist in the CFG and thus must then be enumerated directly, making it effectively impossible to learn about the exploit behavior from this type of evidence (this precludes any analysis of exploit root causes, as discussed below).
@@
4 SABRE
SABRE showcases $\mathcal{C}F_{Log}$-based root cause analysis and remediation with verbatim control flow traces and without the source-code of the attested program. Fig. 3 shows SABRE workflow consisting of five internal modules: Static Analysis, Path Verifier, Exploit Locator, Patch Generator, and Patch Validator. SABRE takes as input an application binary (App.elf) and a verbatim $\mathcal{C}F_{Log}$ (E2 evidence type). The notation used in the rest of the paper is summarized in Tab. I.
Table I: Notation Summary
@@ 4 SABRE
$loop_{count}$ a special entry in $\mathcal{C}F_{Log}$ to optimize static loops (i.e., loops with no internal branches) used by prior works [49, 6]; it denotes the total times the previous $\mathcal{C}F_{Log}$ entry repeated
$addr_{target}$ The control data (e.g., return address, indirect jump target, or function pointer) that was corrupted
$\mathcal{C}F_{Slice}$ A slice of $\mathcal{C}F_{Log}$ that starts where $addr_{target}$ was initialized and ends where the corrupted $addr_{target}$ was used.
@@
$MemMap$ map between memory addresses and values for memory addresses that are in use during Symbolic Data Flow Analysis.
$RegMap$ map between registers and values for registers that are in use during Symbolic Data Flow Analysis.
$X$ special symbolic value assigned to $base$ during Symbolic Data Flow Analysis to monitor memory accesses to it.
@@
4.1 Static Analysis
The first module of SABRE performs static analysis of the application binary (App.elf) to produce the CFG, which is used as input for subsequent modules. The CFG comprises nodes, where each node represents a set of contiguous (non-branching) instructions starting after the previous branch instruction and ending with the next branch, as shown in Fig. 4. This module is executed statically, once per App.elf. The remaining modules are executed for each received $\mathcal{C}F_{Log}$.
Figure 4: Example CFG and $\mathcal{C}F_{Log}$
4.2 Path Verifier
SABRE interprets each $\mathcal{C}F_{Log}$ entry as either: (1) a branch destination (denoted $addr_{Ni}$ in Fig. 4) or (2) a loop counter (denoted $loop_{count}$ in Fig. 4), indicating the number of times that a CFG node in the previous entry has executed. After receiving $\mathcal{C}F_{Log}$, Path Verifier uses this information to determine whether $\mathcal{C}F_{Log}$ constitutes a valid control flow path by traversing the CFG and SS, as described in Sec. 3.2. Upon detecting an invalid transfer, it proceeds to analyze the type of control flow instruction (i.e., ret or indirect call) whose destination address is being corrupted. This is done by first locating the node in the CFG that aligns with the last valid transfer in $\mathcal{C}F_{Log}$. The corrupted instruction is assigned as the last instruction of this CFG node. Using $App.elf$, the corrupted instruction and target address ($addr_{target}$) can be identified as either a return address or an indirect branch target. As a result, this module outputs the corrupted instruction and $addr_{target}$ to the next module of SABRE.
4.3 Exploit Locator
This module pinpoints exploit sources in three phases.
4.3.1 Phase 1. Backwards Traversal
Given the corrupted instruction from the Path Verifier, this phase attempts to find:
@@ 4.3 Exploit Locator
    $\mathcal{C}F_{Slice}$: the section of $\mathcal{C}F_{Log}$ relevant to this particular exploit, which includes the $\mathcal{C}F_{Slice}$ entries between the time when memory storing $addr_{target}$ was assigned with the last pre-corruption value and the time that the corrupted branch executes causing the detected violation.
@@
The procedure for determining $\mathcal{C}F_{Slice}$ and $base$ varies based on the type of $addr_{target}$. This is shown in Fig. 5 and Fig. 6 and described below:
Figure 5: Backward Traversal for a Return Address
Figure 6: Backward Traversal for an Indirect Call
  • $\bullet$
    Corrupted returns. Return addresses are initialized when pushed onto the stack during function calls, and they remain unmodified until the respective returns. Consequently, any corruption to a return address must have occurred between the function call and the return instruction. Therefore as a first step, App.elf is inspected to determine the start address of the function containing the corrupted return. In the example from Fig. 5, Path Verifier detects that the 17th $\mathcal{C}F_{Log}$ entry is invalid, corresponding to the corrupted return instruction at address 0xe152 and $addr_{target}$ pointing to address 0xf078. This module then inspects the disassembled instructions of App.elf and identifies that function F2 contains this corrupted instruction. Then, it saves the start address of F2, 0xe0b6, as the target of the backward traversal. A backward traversal of $\mathcal{C}F_{Log}$ is performed to find the last call to F2, which happens to be at the 8th $\mathcal{C}F_{Log}$ entry. As a result, this phase returns $\mathcal{C}F_{Slice}$ consisting of entries 8 to 17, i.e., $\mathcal{C}F_{Log}$[8:17]. In addition, since return addresses are always initialized on the stack, $base$ can be found in the stack pointer register (sp) at the beginning of $\mathcal{C}F_{Slice}$. Hence, this phase also outputs $base$ to be sp.
  • $\bullet$
@@
4.3.2 Phase 2. Symbolic Data Flow Analysis
@@
$base$ : symbol holding $addr_{target}$ at $\mathcal{C}F_{Slice}$
start;
@@
$MemMap$ : memory map containing
address-value pairs used in $\mathcal{C}F_{Slice}$;
$RegMap$ : register map containing
register-value pairs used in $\mathcal{C}F_{Slice}$;
$FreeList$ : list of freed pointers during
evaluation of $\mathcal{C}F_{Slice}$;
@@
Within $\mathcal{C}F_{Slice}$ determined by Phase 1, Exploit Locator must locate the instruction that corrupts $addr_{target}$, typically requiring emulation of $\mathcal{C}F_{Slice}$ execution. Without access to run-time data prior to $\mathcal{C}F_{Slice}$ execution, Exploit Locator emulates this execution using symbolic analysis. Unlike traditional symbolic execution, we do not use symbolic analysis to explore different paths. Instead, we use it to reconstruct all possible data flows during $\mathcal{C}F_{Slice}$ execution [14, 15] and monitor for the change in data flows that result in corruption of $addr_{target}$.
  • $\bullet$
  • $\bullet$
  • $\bullet$
The concrete value of base is only available during run-time and cannot be determined from $\mathcal{C}F_{Slice}$. Therefore, this phase initializes base with a special symbolic value ($X$) and updates the state (either MemMap or RegMap) in which base is stored accordingly. For instance, in Fig. 5 and 6, as base can be found in sp at the beginning of $\mathcal{C}F_{Slice}$, it sets $Reg$[sp] to $X$.
@@
Next, Exploit Locator uses the evaluation results to determine whether a memory write to $base$ has occurred in the current evaluation. This is done by checking if a new value has been inserted to address $X$ in the updated $MemMap$. If so, it indicates that the current instruction introduces a value (i.e., writes) to memory at $base$, leading to $addr_{target}$’s corruption. Exploit Locator then outputs the address of this instruction as $addr_{Acc}$, along with $MemMap$, $RegMap$, and FreeList to the next phase.
When no write to $X$ is detected, it continues with symbolic execution over the remaining $\mathcal{C}F_{Slice}$ entries. We also note that this analysis may encounter a $\mathcal{C}F_{Slice}$ entry in the form of $loop_{count}$ (as depicted in Fig. 4). In this case, it repeats symbolic evaluation over the same CFG node for $loop_{count}$ times.
4.3.3 Phase 3. Detecting Exploit Type
In its final phase, the Exploit Locator attempts to determine whether the control flow attack resulted from a use-after-free or a buffer overflow, using the output from the previous phase.
Use-after-free. The FreeList can be inspected to determine whether a use-after-free has occurred. Since the previous phase stopped at the point where base was overwritten, the Exploit Locator checks the contents of FreeList to see if it contains base. If base is in FreeList and it was illegally modified after it was freed, it is likely that base was corrupted due to a use-after-free. In this case, Exploit Locator outputs the candidate attack root cause as a use-after-free.
Buffer Overflows. The Exploit Locator assumes that an attack exploits a buffer overflow when base is not in FreeList and when the CFG node containing $addr_{acc}$ has been executed multiple times in $\mathcal{C}F_{Slice}$ (implying that this CFG node is part of a loop that performs repeated memory writes, eventually overflowing into the bounds of $addr_{target}$). In this case, Exploit Locator outputs the attack root cause as a buffer overflow.
Figure 7: Backward traversal for determining buffer bounds.
Figure 8: Example patch for use-after-free vulnerabilities
4.4 Patch Generator

### v2

@@ 3.1. Evidence Types
Figure 2. A conditional branch (CB), indirect/direct calls (IC/DC), and return (RET) in (a) hash-based, (b) verbatim, and (c) hybrid evidence.
@@
Hash-based Evidence (E1). As control flow transfers occur, branch destination addresses are accumulated into a single hash chain. This is performed by first accumulating the current destination address with the previous hash and then hashing the accumulated result (Abera et al., 2016; Dessouky et al., 2017; Huo et al., 2020; Zeitouni et al., 2017), as exemplified in Fig. 2.a. After the first branch occurs, its destination 0xe004 is accumulated into the hash. Since this is the first branch destination, it is accumulated with an initial constant to obtain $H_{0}$. Upon the second branch instruction, its destination address 0xf000 is accumulated with $H_{0}$ to obtain $H_{1}$. This process continues until the end of the requested execution generating a unique value that represents the path. Although simple control flows produce a single hash in E1, programs with more complex control flows (including those with nested loops and branching) might require a set of hash digests (Abera et al., 2016; Zeitouni et al., 2017). This approach yields minimal storage/transmission overheads (i.e., the fixed size of a hash digest). However, $\sf{\mathcal{V}rf}$ must perform its analysis solely based on the final hash digest and the executable binary.
@@
Verbatim Evidence (E2). On the other side of the spectrum, more recent CFA techniques advocate for storing and transmitting $\mathcal{C}F_{Log}$ in its entirety (Nunes et al., 2021b; Dessouky et al., 2018; Toffalini et al., 2019; Caulfield et al., 2023; Zhang et al., 2021; Caulfield et al., 2024a). As depicted in Fig. 2.b, the full sequence of destination addresses is maintained in a buffer. Therefore, this approach requires further (lossless) compression or is limited to small and self-contained operations to avoid filling the dedicated storage too quickly. A common strategy is to log simple loops (those without internal branches) with their address once and a count/condition pertaining to the number of iterations performed (Caulfield et al., 2023; Nunes et al., 2021b; Zhang et al., 2021; Caulfield et al., 2024a). Another reduction is obtained by recording statically defined branch destinations as a single bit (Dessouky et al., 2018). Since static destinations can be determined from the binary alone, branch instructions with statically defined destinations can be recorded as a single bit (1/0) to determine if the branch was taken or not. Branches with dynamically defined destinations (e.g., returns and indirect calls) are still recorded with their full addresses. If $\mathcal{C}F_{Log}$ remains larger than available storage despite these optimizations, it can be split into a set of multiple fixed-size slices and streamed to $\sf{\mathcal{V}rf}$, as suggested in (Toffalini et al., 2019; Caulfield et al., 2023, 2024a). Despite the aforementioned challenges, E2 provides $\sf{\mathcal{V}rf}$ with full control flow path information for analysis.
@@
Hybrid Evidence (E3). Some techniques implement a combination of E1 and E2. OAT (Sun et al., 2020) and ARI (Wang et al., 2023) record forward edges verbatim and accumulate backward edges (returns) into a single hash chain. As shown in Fig. 2.c, the RoT in $\sf{\mathcal{P}rv}$ maintains a trace of forward edges ($T_{F}$) and a hash chain of returns ($H_{R}$). Since 0xf000 is the destination of an indirect call, its full address must be recorded whereas the first destination can be represented by a single bit ‘0’ encoding the ‘not-taken’ conditional branch address. After the attested execution ends, $T_{F}$ and $H_{R}$ are sent to $\sf{\mathcal{V}rf}$.
3.2. Detection & Remediation via E1, E2 & E3
@@
Consider an attack detection strategy in which $\sf{\mathcal{V}rf}$ constructs the CFG of the executable binary, i.e., as suggested in (Abera et al., 2016; Caulfield et al., 2023; Dessouky et al., 2017; Zeitouni et al., 2017; Toffalini et al., 2019; Huo et al., 2020; Zhang et al., 2021). An initial set of invalid paths is defined as forward edges that do not exist in the statically constructed CFG. As valid indirect targets (e.g., return addresses or indirect calls) are over-approximated using the static CFG (Burow et al., 2019), $\sf{\mathcal{V}rf}$ may also use $\mathcal{C}F_{Log}$ along with the program’s binary to emulate the execution of the program locally. In this process, a Shadow Stack (SS) can be emulated locally to verify backward edges (return destinations) and data flow (DF) analysis can be employed to determine a set of valid call sites for each indirect call, updating the CFG accordingly. We observe feasibility of this verification strategy varies greatly depending on evidence types E1, E2, and E3.
Figure 3. SABRE submodules and internal workflow
@@
For E1, with the only evidence available being a hash digest, $\sf{\mathcal{V}rf}$ must infer which path has resulted in the digest. To that end, it must generate each possible legal path based on the CFG and use an emulated SS to accumulate the path’s hash during each step of the CFG traversal. However, the complexity of this task grows exponentially with the number of branches, leading to the intractable path explosion problem (Ramalingam, 1994; Baldoni et al., 2018). Considering anomalous/illegal paths, the analysis is even more challenging. $\sf{\mathcal{V}rf}$ would need to map the hash digest to the exact illegal path. Illegal paths do not exist in the CFG and thus must then be enumerated directly, making it effectively impossible to learn about the exploit behavior from this type of evidence (this precludes any analysis of exploit root causes, as discussed below).
@@
4. SABRE
SABRE showcases $\mathcal{C}F_{Log}$-based root cause analysis and remediation with verbatim control flow traces (E2) and without the source-code of the attested program. Fig. 3 shows SABRE workflow consisting of five internal modules: Static Analysis, Path Verifier, Exploit Locator, Patch Generator, and Patch Validator. SABRE takes as input an application binary (App.elf) and a verbatim $\mathcal{C}F_{Log}$ (E2 evidence type). The notation used in the rest of the paper is summarized in Table 1.
Table 1. Notation Summary
@@ 4. SABRE
$loop_{count}$ a special entry in $\mathcal{C}F_{Log}$ to optimize static loops (i.e., loops with no internal branches) (Nunes et al., 2021b; Caulfield et al., 2023); denotes total times the previous $\mathcal{C}F_{Log}$ entry repeated
$addr_{target}$ The control data (e.g., return address, indirect jump/call target) that was corrupted
$\mathcal{C}F_{Slice}$ A $\mathcal{C}F_{Log}$ slice from $addr_{target}$’s initialization point its corrupted usage.
@@
$MemMap$ address-value mapping for in-use memory addresses during Symbolic DF Analysis.
$RegMap$ register-value mapping for in-use registers during Symbolic DF Analysis.
$FreeList$ list of all freed pointers during Symbolic DF Analysis
$X$ special symbolic value assigned to $base$ during Symbolic DF Analysis.
@@
4.1. Static Analysis and Path Verifier
The first two stages of SABRE are well studied from prior works (Ammar et al., 2024). The first module of SABRE performs static analysis of the application binary (App.elf) to produce the CFG used by subsequent modules. Each CFG node represents a set of contiguous (non-branching) instructions starting with the previous branch destination and ending with the next branch, as shown in Fig. 6. This module is executed once per App.elf. The remaining modules are executed for each received $\mathcal{C}F_{Log}$.
Figure 4. CFG and $\mathcal{C}F_{Log}$
Figure 5. Backward Traversal for a ret
Figure 6. Backward Traversal for an call
SABRE verifies the contents of the verbatim $\mathcal{C}F_{Log}$, the contents of which are assumed to have the format depicted in Fig. 6. SABRE interprets each $\mathcal{C}F_{Log}$ entry as either: (1) a branch destination (denoted $addr_{Ni}$ in Fig. 6) or (2) a loop counter (denoted $loop_{count}$ in Fig. 6), indicating the number of times that a CFG node in the previous entry has executed. After receiving $\mathcal{C}F_{Log}$, Path Verifier uses this information to determine whether $\mathcal{C}F_{Log}$ constitutes a valid control flow path by traversing the CFG and SS, as described in Sec. 3.2. Upon detecting an invalid transfer, it proceeds to analyze the type of control flow instruction (i.e., ret or indirect call) whose destination address is being corrupted. This is done by first locating the node in the CFG that aligns with the last valid transfer in $\mathcal{C}F_{Log}$. The corrupted instruction is assigned as the last instruction of this CFG node. Using $App.elf$, the corrupted instruction and target address ($addr_{target}$) are identified as either a return address or an indirect branch target. This module outputs the corrupted instruction and $addr_{target}$ to the next module of SABRE.
4.2. Exploit Locator
Phase 1. Backwards Traversal. Given the corrupted instruction from the Path Verifier, this phase attempts to find:
@@ 4.2. Exploit Locator
    $\mathcal{C}F_{Slice}$: the section of $\mathcal{C}F_{Log}$ relevant to this particular exploit, including the $\mathcal{C}F_{Slice}$ entries between the time when memory storing $addr_{target}$ was assigned with the last pre-corruption value and the time that the corrupted branch executes causing the detected violation.
@@
The procedure for determining $\mathcal{C}F_{Slice}$ and $base$ varies based on the type of $addr_{target}$. This is shown in Fig. 6 and Fig. 6.
Corrupted returns. Return addresses are initialized when pushed onto the stack during function calls, and they remain unmodified until the respective returns. Consequently, any corruption to a return address must have occurred between the function call and the return instruction. Therefore as a first step, App.elf is inspected to determine the start address of the function containing the corrupted return. In the example from Fig. 6, Path Verifier detects that the 17th $\mathcal{C}F_{Log}$ entry is invalid, corresponding to the corrupted return instruction at address 0xe152 and $addr_{target}$ pointing to 0xf078. This module then inspects the disassembled instructions of App.elf and identifies that function F2 contains this corrupted instruction. Then, it saves the start address of F2, 0xe0b6, as the target of the backward traversal. A backward traversal of $\mathcal{C}F_{Log}$ is performed to find the last call to F2, which happens to be at the 8th $\mathcal{C}F_{Log}$ entry. As a result, this phase returns $\mathcal{C}F_{Slice}$ consisting of entries 8 to 17, i.e., $\mathcal{C}F_{Log}$[8:17]. In addition, since return addresses are always initialized on the stack, $base$ can be found in the stack pointer register (sp) at the beginning of $\mathcal{C}F_{Slice}$. Hence, this phase also outputs $base$ to be sp.
@@
Phase 2. Symbolic DF Analysis. Within $\mathcal{C}F_{Slice}$ determined by Phase 1, Exploit Locator must locate the instruction that corrupts $addr_{target}$, typically requiring emulation of $\mathcal{C}F_{Slice}$ execution. Without access to run-time data prior to $\mathcal{C}F_{Slice}$ execution, Exploit Locator emulates this execution using symbolic analysis. Unlike traditional symbolic execution (King, 1976), we do not use symbolic analysis to explore different paths. Instead, we use it to reconstruct all possible data flows during $\mathcal{C}F_{Slice}$ execution (Yagemann et al., 2021b, a) and monitor for the change in data flows that result in corruption of $addr_{target}$.
  • $\bullet$
  • $\bullet$
  • $\bullet$
The concrete value of base is only available during run-time and cannot be determined from $\mathcal{C}F_{Slice}$. Therefore, this phase initializes base with a special symbolic value ($X$) and updates the state (MemMap or RegMap) in which base is stored accordingly. For instance, in Fig. 6 and 6, as base is found in sp at the start of $\mathcal{C}F_{Slice}$, $Reg$[sp] is set to $X$.
@@
$base$ : symbol holding $addr_{target}$ at $\mathcal{C}F_{Slice}$ start;
@@
$MemMap$ : memory map containing address-value pairs used in $\mathcal{C}F_{Slice}$;
$RegMap$ : register map containing register-value pairs used in $\mathcal{C}F_{Slice}$;
$FreeList$ : list of freed pointers during evaluation of $\mathcal{C}F_{Slice}$;
@@
Next, Exploit Locator uses the evaluation results to determine whether a memory write to $base$ has occurred in the current evaluation. This is done by checking if the value of $X$ in $MemMap$ has changed. If so, it indicates that the current instruction introduces a value (i.e., writes) to memory at $base$, leading to $addr_{target}$’s corruption. Exploit Locator then outputs the address of this instruction as $addr_{Acc}$, along with $MemMap$, $RegMap$, and FreeList to the next phase.
When no write to $X$ is detected, it continues with symbolic execution over the remaining $\mathcal{C}F_{Slice}$ entries. We note that this analysis may encounter a $\mathcal{C}F_{Slice}$ entry that is a $loop_{count}$ (as depicted in Fig. 6). In this case, it repeats symbolic evaluation on the same CFG node for $loop_{count}$ times.
Phase 3. Detecting Exploit Type. In its final phase, the Exploit Locator attempts to determine whether the control flow attack resulted from a use-after-free or a buffer overflow, using the output from the previous phase.
The FreeList can be inspected to determine whether a use-after-free has occurred. Since the previous phase stopped at the point where base was overwritten, the Exploit Locator checks the contents of FreeList to see if it contains base. If base is in FreeList and it was illegally modified after it was freed, it is likely that base was corrupted due to a use-after-free. In this case, Exploit Locator outputs the candidate attack root cause as a use-after-free.
The Exploit Locator assumes that an attack exploits a buffer overflow when base is not in FreeList and when the CFG node containing $addr_{acc}$ has been executed multiple times in $\mathcal{C}F_{Slice}$ (implying that this CFG node is part of a loop that performs repeated memory writes, eventually overflowing into $addr_{target}$). In this case, Exploit Locator outputs the attack root cause as a buffer overflow.
Figure 7. Example patch for use-after-free vulnerabilities
Figure 8. Backward traversal for determining buffer bounds.
4.3. Patch Generator